Find the distances between the following pairs of parallel lines
(a) r · (i + j) + 7 = 0, r · (i + j) − 11 = 0;
(b) r · (2i − 3j) + 6 = 0, r · (4i − 6j) + 5 = 0;
(c) r = i + j + s(4i − j), r = 4i + 5j + t(8i − 2j).
1
Expert's answer
2019-11-01T13:18:53-0400

(a)

converting equation into general form

then given lines are-


"x+y+7=0\\and\\x+y-11=0\\distance=\frac{|d_1-d_2|}{\sqrt{a^2+b^2}}=\frac{|7+11|}{\sqrt{1^2+1^2}}=\frac{18}{\sqrt{2}}=9\sqrt{2}=12.728"

(b)

Similarly given lines are -


"4x-6y+12=0\\and\\4x-6y+5=0\\distance=\frac{|d_1-d_2|}{\sqrt{a^2+b^2}}=\frac{|12-5|}{\sqrt{4^2+(-6)^2}}=\frac{7}{\sqrt{52}}=0.97"

(c)

passing points are "(1,1)\ and\ (4,5)"

distance betwwen passing point of these two lines is-


"d=\sqrt{(x_1-x_2)^2+(y_1-y_2)^2}=\sqrt{(4-1)^2+(5-1)^2}\\d=5"

vector between these two points= A = "3\hat i+4\hat j"

vector parallel to any of the two line =B = "4\hat i-\hat j"

Distance between two lines = "d\sin\theta"

where "\theta" is the angle between vector A and B

"now,\\\cos\theta=\frac{\overset{\to}A.\overset{\to}B}{|\overset{\to}A||\overset{\to}B|}=\frac{12-4}{5\times\sqrt{17}}=\frac{8}{5\sqrt{17}}\\then,\sin\theta=0.922"

So,

distance between two lines = "d\sin\theta=5\times0.922=4.61"